During March 2019, Alaska Corporation recorded $269,000 of costs related to factory overhead. Alaska's overhead application rate is based on direct labor hours. The preset formula for overhead application estimated that $250,000 would be incurred, and 12,500 direct labor hours would be worked. During March, 12,400 hours were actually worked. Use this information to determine the amount of overhead over or under applied. Enter overapplied overhead as a negative number. (round & enter any final dollar answers to the nearest whole dollar
Answer #1

Predetermined overhead rate=Estimated overhead/Estimated direct labor hours

=(250,000/12500)=$20 per direct labor hour

Applied overhead=Predetermined overhead rate*Actual direct labor hours

=(20*12400)=$248000

Hence since applied overhead is less than actual overhead;

Under applied overhead=269,000-248000

=$21000